OpenStudy (anonymous):

supplementary angles. m is 4y + 7 and m is 9y + 4. Find m. (2 points)

OpenStudy (anonymous):

I don't see what this has to do with supplementary angles..... Anyway, m = 4y +7 = 9y + 4 gives you y which then gives you m

OpenStudy (anonymous):

<1 and <2 are supplementary angles. M<1 is 4y+7 and M<2 = 9y+4. find M<2 thats what it really says it didnt show sorry

OpenStudy (anonymous):

The answer is more or less same, anyway.... 4y +7 = 9y + 4 gives you y which then gives you (either angle)